### v1

# Fitting regular point patterns with a hyperuniform perturbed lattice

Abstract
We introduce a new methodology for modeling regular spatial data using hyperuniform point processes. We show that, under some mixing conditions on the perturbations, perturbed lattices in general dimension are hyperuniform. Due to their inherent repulsive structure, they serve as an effective baseline model for data sets in which points exhibit repulsiveness. Specifically, we derive an explicit formula for the $K$-function of lattices perturbed by a Gaussian random field, which proves particularly useful in conjunction with the minimal contrast method. We apply this approach to a data set representing the grain centers of a polycrystalline metallic material composed of nickel and titanium.
Keywords. hyperuniformity, perturbed lattices, mixing, random fields, minimal contrast, polycrystalline microstructure
AMS 2010 Subject Classification: Primary: 60G55; Secondary: 60D05, 60K35
1 Introduction
Hyperuniformity, often described as a state of global order and local disorder, refers to a unique characteristic of certain point processes (or structures in general) whose fluctuation in number of points is suppressed compared to the Poisson point process. This property appears in many scientific fields. For instance, the eigenvalues of Gaussian random matrices were shown to be hyperuniform, as well as the zeros of Gaussian random polynomials. In biology, avian photoreceptor patterns [9] exhibit hyperuniformity, as do receptors in the immune system [17]. In particular, the concept of hyperuniformity is relevant in the study of physical systems, offering insights into the organization of particles at large scales despite their seemingly disordered local appearance (e.g. the order of the early Universe [19]). For a more in-depth discussion, see [23] for a physical perspective and [2] for a mathematical treatment.
In many cases, regularity comes hand in hand with hyperuniformity of the process. However, it is not a rule, as we know of examples of hyperfluctuating regular point processes. Regular patterns appear in many natural contexts, such as in cellular structures and crystals of different materials. In material science, repulsive phenomena are often modeled using Gibbs models with repulsive energy functions. Although such models generally fit observed data well, fluctuations in the number of points per unit volume are Poisson-like, as noted in [5].
However, if there is evidence of hyperuniformity, either from the physical nature of the data or from a given statistical test (see recent results [12, 8, 16]), then important questions arise. What model should we choose and how to efficiently generate large samples from such a model for further statistical inference? While some determinantal point processes, sine-$\beta$ processes or Coulomb gases are known to be hyperuniform, generating large samples of these processes remains computationally challenging. Until recently, the only way to limit the computation times was to start with a fixed hyperuniform structure such as $\mathbb{Z}^{d}$ and displace all the points independently. Unfortunately, this model preserves very strong underlying bindings (this is related to Bragg peaks) emerging from the original structure. Although such a model is computationally advantageous, this flaw makes the model too restrictive as seen in our analysis of the data of nitinol centers in Section 4.
A promising new alternative lies in dependent displacements. It was shown in [4] that the finite second moment of the common law of perturbations ensures that the resulting process is hyperuniform in dimensions $d=1,2$ regardless of the covariance structure among the perturbations. However, in dimensions $d\geq 3$, this is far from true. Even almost surely bounded perturbations of the higher dimensional lattice can produce non-hyperuniform structures.
There are other theoretical arguments supporting the choice of a hyperuniform perturbed lattice as a ground model for our data. From a statistical standpoint, reducing fluctuation order in such models could enhance computational efficiency in numerical integrations. The slower the number variance increases, the better are the new samples to estimate statistics of the process $\mu$ of the type $\int f\mathrm{d}\,\mu$. Moreover, recent results in [13] demonstrate that under certain integrability conditions on the pair correlation function of the process, any hyperuniform point process in $d\leq 2$ lies within a finite Wasserstein distance from an $L_{2}$-perturbed lattice. Given these facts, since perturbed lattices naturally preserve the regularity of the underlying lattice, it makes them an attractive choice for fitting experimental regular data.
In the present paper, we provide options for modeling regular point pattern by hyperuniform perturbed lattices in general dimension. We formulate conditions for hyperuniformity that take into account not only moments but also the covariance structure via the $\alpha$-mixing coefficient. If $\mathbb{E}|p_{0}|^{r}<\infty$ for some $r>1$, then
$\sum_{i\in\mathbb{Z}^{d}}\alpha^{\frac{r}{r-1}}(|i|)<\infty$ (1.1)
is sufficient for hyperuniformity of the perturbed lattice
$\Xi=\{i+p_{i}+U;i\in\mathbb{Z}^{d}\}.$
Specially, Gaussian perturbations with fast decaying correlations satisfy (1.1) and, moreover, it is possible to find and explicit formula for the respective $K$-function. That makes the model fitting even faster and more straightforward. The results in this paper are accompanied by a study of real spatial data that were reconstructed from an experimental 3D synchrotron x-ray diffraction (3D-XRD). It is demonstrated in Section 4 that choosing a suitable correlation structure is a crucial and delicate task.
2 Main results
Assume that $\mathcal{L}$ is some lattice in $\mathbb{R}^{d}$ (e.g. $\mathbb{Z}^{d}$ or its affine transformation). Perturbed lattices, i.e. stationary point processes of the form $\Xi=\{i+p_{i}+U,i\in\mathcal{L}\}$, are nowadays receiving a lot of interest ([7, 4, 2, 11]). Until recently, independently perturbed lattices in general dimension and dependently perturbed lattices in $d=1,2$ were considered important representatives among hyperuniform point processes. Here we state conditions on the $\alpha-$mixing coefficient of the perturbation field $\mathbf{p}=\{p_{i},i\in\mathcal{L}\}$ under which the perturbed lattice $\Xi$ preserves hyperuniformity in higher dimensions as well. The proofs are postponed to Section 2.5 after building up the necessary tools.
Theorem 1.
Let $\mathbf{p}=\{p_{i},i\in\mathcal{L}\}$ be a strictly stationary random field such that there exist $s,r>0$ with $11/s+1/r=$ 1, $\mathbb{E}|p_{0}|^{r}<\infty$ and
$\sum_{i\in\mathcal{L}}\alpha^{1/s}(|i|)<\infty.$
Then the perturbed lattice $\Xi=\mathcal{L}+\mathbf{p}+U$ is hyperuniform.
Theorem 2.
Let $\mathbf{p}=\{p_{i},i\in\mathcal{L}\}$ be a strictly stationary random field with almost surely bounded perturbations. If
$\sum_{i\in\mathcal{L}}\alpha(|i|)<\infty,$
then the perturbed lattice $\Xi=\mathcal{L}+\mathbf{p}+U$ is hyperuniform.
Corollary 1.
For some $m\in\mathbb{N}$, let $\mathbf{p}=\{p_{i},i\in\mathcal{L}\}$ be a strictly stationary $m$-dependent random field such that $\mathbb{E}|p_{0}|^{\delta}<\infty$ for some $\delta>1$. Then the perturbed lattice $\Xi=\mathcal{L}+\mathbf{p}+U$ is hyperuniform.
Corollary 2.
Let $\mathbf{p}=\{p_{i},i\in\mathbb{Z}^{d}\}$ be a strictly stationary Gaussian random field with $\mathrm{cov}(p_{0},p_{i})|=O(|i|^{-\gamma})$ for some $\gamma>2d$ and positive spectral density. Then the perturbed lattice $\Xi=\mathbb{Z}^{d}+\mathbf{p}+U$ is hyperuniform.
2.1 General notation
To describe spatial point processes, i.e. point processes on $\mathbb{R}^{d}$, $d\geq 1$, we denote by $\mathcal{B}^{d}$ be the Borel $\sigma-$algebra on $\mathbb{R}^{d}$ and $\mathcal{B}_{b}^{d}$ the system of all bounded Borel sets. Moreover, $\lambda_{d}$ denotes the Lebesgue measure on $(\mathbb{R}^{d},\mathcal{B})$, ||$\cdot$|| ⋅ | is the Euclidean norm, and #$\#(A)$ stands for the cardinality of a set $A$. For the purpose of counting points in balls, we denote by $B(x,r)$ the closed Euclidean ball centered in $x$ with radius $r>0$. In particular, we use the abbreviation $B_{r}$ when $x=0$. If $x,y\in\mathbb{R}^{d}$ are two $d-$dimensional vectors, we denote by $x\cdot y$ the usual dot product in $\mathbb{R}^{d}$.
In addition, we assume that all random variables are defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$, which remains fixed throughout the whole paper. If two random variables $X,Y$ are equal in distribution, we write $X\stackrel{{\scriptstyle D}}{{=}}Y$.
2.2 Point processes and second order measures
At this place, we recall some terminology related to point processes. As usual, let $\mathbf{N}$ be the space of all locally finite point configurations of $\mathbb{R}^{d}$, i.e.
$\mathbf{N}:=\left\{\mu\subset\mathbb{R}^{d};\#(\mu\cap B)<\infty,\,\forall B\in\mathcal{B}^{d}_{b}\right\}.$
Further, $\mathcal{N}$ is the smallest $\sigma$-algebra on $\mathbf{N}$ making the mapping $\mu\mapsto\#(\mu\cap B)$ measurable for each $B\in\mathcal{B}^{d}$. Specifically, $\mathbf{N}_{f}$ is a subset of $\mathbf{N}$ consisting of only finite configurations.
A point process is a random element $\mu$ defined on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $(\mathbf{N},\mathcal{N})$. We call a point process $\mu$ stationary if its distribution is invariant with respect to shifts, meaning that $\mu\stackrel{{\scriptstyle D}}{{=}}\mu+x$ for all $x\in\mathbb{R}^{d}$.
Definition 1 (m-th factorial moment measure).
Let $m\in\mathbb{N}$. For a point process $\mu$, we define the $m$-th factorial moment measure as the Borel measure $\alpha^{(m)}$ on $(\mathbb{R}^{d})^{m}$ for which
$\alpha^{(m)}(B_{1}\times\cdots\times B_{m}):=\mathbb{E}\sum_{x_{1},\ldots,x_{m}\in\mu}^{\neq}\mathbf{1}(x_{1}\in B_{1},\ldots,x_{m}\in B_{m}),$
where $B_{1},\ldots,B_{m}\in\mathcal{B}^{d}.$
Specifically, we call $\alpha^{(1)}$ the intensity measure. If $\mu$ is stationary, then it is simply a multiple of the Lebesgue measure, i.e. $\alpha^{(1)}=\rho\lambda_{d}$ for some $\rho\geq 0$. Moreover, desintegrating $\alpha^{(2)}$ of a stationary point process leads to the definition of reduced second factorial moment measure $\alpha^{(2)}_{red}$ rather known as pair correlation measure determined by a relation
$\alpha_{red}^{(2)}(B)=\mathbb{E}\sum_{x,y\in\mu}^{\neq}\mathbf{1}\{x\in[0,1]^{d},y-x\in B\},\quad B\in\mathcal{B}^{d}.$ (2.1)
If, in addition, $\alpha^{(2)}_{red}$ is absolutely continuous w.r.t. $\lambda_{d}$, then the respective density $g$ is called the pair correlation function.
Definition 2 ($m$-th factorial cumulant measure).
For $m\in\mathbb{N}$, we define the $m$-th factorial cumulant measure of a point process $\mu$ as a measure on $(\mathbb{R}^{d})^{m}$ by
$\gamma^{(m)}(B_{1}\times\cdots\times B_{m})=\sum_{j=1}^{m}(-1)^{j-1}(j-1)!\sum_{K_{1}\cup\cdots\cup K_{j}=\{1,\ldots,m\}}\prod_{r=1}^{j}\alpha^{(\#(K_{r}))}(\times_{s\in K_{r}}B_{s}).$
Factorial cumulant measures form a useful tool in expressing and studying dependencies among distant parts of a point pattern. The relation between factorial moment measures and factorial cumulant measures is based on the general relationship between mixed moments and mixed cumulants. The first factorial cumulant measure coincides with the intensity measure (first moment measure), whereas the second factorial cumulant measure equals to
$\gamma^{(2)}(B_{1}\times B_{2})$ $=\alpha^{(2)}(B_{1}\times B_{2})-\alpha^{(1)}(B_{1})\alpha^{(1)}(B_{2})$
$=\textup{cov}(\#(\mu\cap B_{1}),\#(\mu\cap B_{2}))-\alpha^{(1)}(B_{1}\cap B_{2}).$
The second line is the reason why $\gamma^{(2)}$ (or its normalization) is sometimes referred to as covariance measure. If $\mu$ is stationary, then $\gamma^{(2)}$ (as well as all above mentioned measures) can be disintegrated so that
$\gamma^{(2)}(A\times B)=\rho\int_{A}\gamma^{(2)}_{red}(B-x)\mathrm{d}\,x$
and $\gamma^{(2)}_{red}$ is a unique measure on $\mathbb{R}^{d}$ called reduced second factorial cumulant measure (short RSFC measure) of $\mu$. The RSFC measure is a signed measure, hence admits the Hahn–Jordan decomposition meaning that there exists uniquely determined measures $\gamma^{(2)}_{red,+}$ and $\gamma^{(2)}_{red,-}$ such that
$\gamma^{(2)}_{red}=\gamma^{(2)}_{red,+}-\gamma^{(2)}_{red,-}.$ (2.2)
The total variation measure of $\gamma^{(2)}_{red}$ is then
$|\gamma^{(2)}_{red}|_{TV}=\gamma^{(2)}_{red,+}+\gamma^{(2)}_{red,-}.$
At last, there is a correspondence between the RSFC measure and the second factorial moment measure
$\gamma_{red}^{(2)}(B)=\alpha_{red}^{(2)}(B)-\rho^{2}\lambda_{d}(B).$ (2.3)
2.3 Hyperuniformity of a point process
Definition 3 (Hyperuniform point process).
A stationary point process $\mu$ is called hyperuniform, if
$\sigma(r):=\frac{\mathrm{var}(\#(\mu\cap B_{r}))}{\lambda_{d}(B_{r})}\xrightarrow{r\to\infty}0.$ (2.4)
We can classify the hyperuniformity by the speed of the convergence of $\sigma$. It was proposed in the seminal paper [23] to call a process $\mu$
  • class-I hyperuniform if $\sigma(r)\sim cr^{-1}$
  • class-II hyperuniform if $\sigma(r)\sim cr^{-1}\log(r)$
  • class-III hyperuniform if $\sigma(r)\sim cr^{-\alpha}$, $\alpha\in(0,1)$.
The papers [23, 2] offer a wide range of examples from mathematical and physical perspective. However, as mentioned in [4], the classification is not exhaustive and does not contain all hyperuniform point processes.
Especially in physics papers, hyperuniformity is understood in terms of Bartlett’s spectral measure, whose density (if it exists) is called the structure factor. Hyperuniformity is then characterized by vanishing of the structure factor at frequency $k=0$. The usual definition goes as follows:
Definition 4 (Structure factor).
If $\mu$ is a stationary point process with $|\gamma^{(2)}_{red}|(\mathbb{R}^{d})<\infty$, then we may define the structure factor of $\mu$ by
$S(k)=1+\int_{\mathbb{R}^{d}}e^{-ik\cdot x}\gamma_{red}^{(2)}(\mathrm{d}\,x),\quad k\in\mathbb{R}^{d}.$
The hyperuniformity of $\mu$ then translates to $S(0)=0$. Under the assumption of Definition 4, we may assume that the structure factor follows a power-law behavior near zero:
$S(k)=O(|k|^{\alpha})$ (2.5)
where $\alpha\geq 0$ is called the hyperuniformity exponent. Hyperuniformity then occurs with $\alpha>0$. There is a correspondence between the hyperuniformity exponent and (2.4). Namely, $\alpha\in(0,1)$ if $\mathrm{var}(\#(\mu\cap B_{r}))=O(r^{d-\alpha})$ and $\alpha>1$ corresponds to $\mathrm{var}(\#(\mu\cap B_{r}))=O(r^{d-1})$, i.e. class-I hyperuniformity.
It is discussed in [7] (Chap VI.) that independently perturbed lattices with finite second moment of the perturbations satisfy (2.5) with $\alpha=2$ meanwhile independent perturbations with tail probabilities of the perturbations $\mathbb{P}(|p_{0}|\geq x)=O(x^{-d-\delta})$ for some $\delta\in(0,2]$ satisfy (2.5) with $\alpha=\delta$. However, we need to be careful with the understanding of the structure factor of perturbed lattices as is can not be generally defined in the sense of Definition 4 since the total variation of $\gamma^{(2)}_{red}$ can be infinite.
2.4 Perturbed lattices
Our core concept is the deterministic lattice $\mathcal{L}$ in $\mathbb{R}^{d}$. The lattice $\mathcal{L}$ is defined as a subgroup of $\mathbb{R}^{d}$ generated by linear combinations with integer coefficients of the vectors of some basis $(\lambda_{1},\ldots,\lambda_{d})$ of $\mathbb{R}^{d}$. An elementary example is the lattice $\mathbb{Z}^{d}$. Any other lattice can be obtained as an affine transformation of $\mathbb{Z}^{d}$. The fundamental domain of $\mathcal{L}$ is denoted by $\mathcal{D}$ and defined by
$\mathcal{D}:=\left\{\sum_{i=1}^{d}t_{i}\lambda_{i},\,\,(t_{1},\ldots,t_{d})\in[0,1)^{d}\right\}.$
For $\mathcal{L}=\mathbb{Z}^{d}$, this is simply $\mathcal{D}=[0,1]^{d}$.
Definition 5 (Perturbed lattice).
Let $\mathbf{p}=\{p_{i},i\in\mathcal{L}\}$ be a strictly stationary random field in $\mathbb{R}^{d}$, i.e.
$\{p_{i},i\in\Lambda\}\stackrel{{\scriptstyle D}}{{=}}\{p_{i+j},i\in\Lambda\}$
for all $\Lambda\subset\mathcal{L}$ and $j\in\mathcal{L}$. Let $U$ be a uniform random variable $U$ on $\mathcal{D}$ independent of $\mathbf{p}$. By a perturbed lattice, we understand a stationary point process $\Xi$ on $\mathbb{R}^{d}$ defined by
$\Xi:=\mathbb{Z}^{d}+U+\mathbf{p}=\cup_{i\in\mathcal{L}}\{i+U+p_{i}\}.$
Specially, if $\mathbf{p}$ consists of independent random variables, we speak about independently perturbed lattice.
By Proposition 2.8 in [4], the reduced factorial moment measure $\alpha^{(2)}$ is given by
$\alpha_{red}^{(2)}(B)=\mathbb{E}\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbf{1}{\{i+p_{i}-p_{0}\in B\}}$ (2.6)
for any Borel $B\subset\mathbb{R}^{d}$. It is now clear that the structure factor of a perturbed lattice can not be defined in the sense of Definition 4 since
$\gamma^{(2)}_{red}(B)=\mathbb{E}\sum_{j\in\mathcal{L}\setminus\{0\}}\mathbf{1}\{j+X_{j}-X_{0}\in B\}-|B|,\quad B\subset\mathbb{R}^{d}\text{ bounded}$
can have infinite total variation.
2.5 Strong mixing of the perturbation field
We recall the notion of the $\alpha$-mixing coefficient that measures the strength of dependence between two $\sigma-$fields. Let $\mathcal{F}_{1},\mathcal{F}_{2}$ be two sub $\sigma$-fields of $\mathcal{F}$. The $\alpha-$mixing coefficient (or strong mixing coefficient) between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ is defined by
$\alpha(\mathcal{F}_{1},\mathcal{F}_{2})=\textrm{sup}\{|\mathbb{P}(A)\mathbb{P}(B)-\mathbb{P}(A\cap B)|;A\in\mathcal{F}_{1},B\in\mathcal{F}_{2}\}.$
For a random variable $X$, we denote $\|X\|_{p}:=(\mathbb{E}|X|^{p})^{1/p}$ for $p<\infty$ and $\|X\|_{\infty}$ the essential supremum of $|X|$ w.r.t. $\mathbb{P}$. If $X$ and $Y$ are two random variables measurable with respect to $\sigma$-fields $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ respectively, then two classical covariance inequalities hold (see Theorem 3 of [6] and its proof):
  1. 1.
    If $\|X\|_{\infty}<\infty$, then for $q,r\geq 1$ such that $q^{-1}+r^{-1}=1$:
    $|\mathrm{cov}(X,Y)|\leq 6\alpha^{1/r}(\mathcal{F}_{1},\mathcal{F}_{2})\|X\|_{\infty}\|Y\|_{q}.$ (2.7)
  2. 2.
    Specially, if $X,Y$ are both a.s. bounded, then
    $|\mathrm{cov}(X,Y)|\leq 4\alpha(\mathcal{F}_{1},\mathcal{F}_{2})\|X\|_{\infty}\|Y\|_{\infty}.$ (2.8)
Let $\mathbf{X}=\{X_{i},i\in\mathcal{L}\}$ be a random field of $\mathbb{R}^{d}$-valued random variables indexed by the lattice $\mathcal{L}$. For a finite subset $C\subset\mathcal{L}$, we denote by $X_{C}=\{X_{i},i\in C\}$ the $C-$marginal distribution of $\mathbf{X}$ and $\sigma(X_{C})$ is the $\sigma$-field generated by $X_{C}$. Moreover, let $d_{\mathcal{L}}$ be some distance on $\mathcal{L}$. Then we define the $\alpha$-mixing coefficient of $\mathbf{X}$ by
$\alpha(n)=\sup_{u,v\in\mathbb{N}}\sup\{\alpha(\sigma(X_{A}),\sigma(X_{B}));A,B\subset\mathcal{L},|A|\leq u,|B|\leq v,d_{\mathcal{L}}(A,B)\geq n\}.$
Finally, we are in a position to prove Theorems 1 and 2. The following lemma will be particularly useful.
Lemma 1.
Let $\mathbf{X}=\{X_{i},i\in\mathcal{L}\}$ be a strictly stationary random field such that $\mathbb{E}|X_{0}|^{r}<\infty$ for some $r>1$. Then there is a constant $c>0$ such that
$\left|\varphi_{X_{i}-X_{0}}(t)-|\varphi_{X_{0}}(t)|^{2}\right|$ $\leq c\min\{\alpha^{1/s}(|i|)|t|\|X_{0}\|_{r},\alpha(|i|)\}\quad\text{for all }t\in\mathbb{R}^{d},$
where $11/s+1/r=$ 1.
Proof (of Theorem 1).
We follow the notation of the proof of Proposition 3.1 of [4]. In fact, the assumption on $\alpha$-mixing coefficient enables drastic simplifications of the respective proof. First, we rewrite the asymptotic variance using Bessel kernels and its Fourier transform. Let $J_{d/2}$ be a Bessel function of the first kind and for $r>0$ we set $j_{r}(t)=C_{d}\frac{J_{d/2}(r|t|)^{2}}{|t|^{d}}$, where $C_{d}$ is chosen so that the Fourier transform of $j_{r}(t)$ satisfies
$\hat{j_{r}}=\frac{\mathbf{1}_{B_{r}}\star\mathbf{1}_{B_{r}}}{|B_{r}|}$
and $\star$ denotes the convolution of two functions. Then the asymptotic variance can be expressed by
$\frac{\mathrm{var}(|\Xi\cap B_{r}|)}{|B_{r}|}=A(r)+B(r),$
with
$A(r)$ $=\sum_{j\in\mathbb{Z}^{d}\setminus\{0\}}\hat{j_{r}}(j)+1-|B_{r}|,$
$B(r)$ $=\mathbb{E}\sum_{j\in\mathbb{Z}^{d}\setminus\{0\}}\hat{j_{r}}(j+X_{j}-X_{0})-\hat{j_{r}}(j).$
Recall that $A(r)=o(1)$ since $A(r)=\mathrm{var}(|\mathbb{Z}^{d}+U\cap B_{r}|)/|B_{r}|$ (asymptotic variance of the hyperuniform stationarized lattice). The residual term can be further expressed using the characteristic function $\varphi_{X_{j}-X_{0}}$ of the difference $X_{j}-X_{0}$:
$B(r)=\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{\mathbb{R}^{d}}e^{-it\cdot j}j_{r}(t)(\varphi_{X_{j}-X_{0}}(t)-1)\mathrm{d}\,t.$
As the mass of $j_{r}$ concentrates around the origin, it is convenient to denote
$B_{1}(r)$ $:=\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{B(0,1)}e^{-it\cdot j}j_{r}(t)(\varphi_{X_{j}-X_{0}}(t)-1)\mathrm{d}\,t$ (2.9)
$B_{2}(r)$ $:=\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{B(0,1)^{C}}e^{-it\cdot j}j_{r}(t)(\varphi_{X_{j}-X_{0}}(t)-1)\mathrm{d}\,t$ (2.10)
so that $B(r)=B_{1}(r)+B_{2}(r).$ In the special case, when the perturbations are i.i.d. having the same distribution as $p_{0}$, we denote the latter quantities by $B^{iid}(r),B^{iid}_{1}(r)$, resp. $B^{iid}_{2}(r)$. We know $B^{iid}(r)\to 0$ as $r\to\infty$ because independently perturbed lattices are known to be hyperuniform.
First, using Lemma 1
$|B_{1}(r)-B^{iid}_{1}(r)|$ $=\left|\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{B(0,1)}e^{it\cdot j}j_{r}(t)\left(\varphi_{p_{i}-p_{0}}(t)-|\varphi_{p_{0}}(t)|^{2}\right)\mathrm{d}\,t\right|$
$\leq\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{B(0,1)}j_{r}(t)\left|\varphi_{p_{i}-p_{0}}(t)-|\varphi_{p_{0}}(t)|^{2}\right|\mathrm{d}\,t$
$\leq c\|p_{0}\|_{r}\int_{B(0,1)}j_{r}(t)|t|\mathrm{d}\,t\sum_{j\in\mathcal{L}\setminus\{0\}}\alpha^{1/s}(|j|).$
This converges to 00 with increasing $r$ since
$\int_{B(0,1)}j_{r}(t)|t|\mathrm{d}\,t\xrightarrow{r\to\infty}0.$
This could be seen by observing that
$0\leq j_{r}(t)|t|\leq j_{r}(t),\quad t\in B(0,1).$ (2.11)
[…]
Similarly, again by Lemma 1,
$|B_{2}(r)-B^{iid}_{2}(r)|$ $\leq\sum_{j\in\mathcal{L}\setminus\{0\}}\int_{B(0,1)^{C}}j_{r}(t)\left|\varphi_{p_{i}-p_{0}}(t)-|\varphi_{p_{0}}(t)|^{2}\right|\mathrm{d}\,t$
$\leq c\int_{B(0,1)^{C}}j_{r}(t)\mathrm{d}\,t\sum_{j\in\mathcal{L}\setminus\{0\}}\alpha(|j|).$
Since $\sum_{j\in\mathcal{L}}\alpha^{1/s}(|j|)$ converges, so does the sum $\sum_{j\in\mathcal{L}\setminus\{0\}}\alpha(|j|).$ An large $|t|$ asymptotic result for the Bessel function (see Section 5.16 of [14]) gives $j_{r}\sim Cr^{-1}x^{-d-1}$ meaning that $\int_{B(0,1)^{C}}j_{r}(t)\mathrm{d}\,t\to 0$ and $|B_{2}(r)-B^{iid}_{2}(r)|\to 0$ with $r\to\infty$ which finishes the proof.
Proof (of Lemma 1).
First, assume that $t\in B(0,1)$. In a similar spirit as in the proof of Theorem 2.1. of [15], we write
$\left|\varphi_{X_{i}-X_{0}}(t)-|\varphi_{X_{0}}(t)|^{2}\right|$ $=\left|\mathbb{E}e^{it\cdot(X_{i}-X_{0})}-\mathbb{E}e^{it\cdot X_{i}}\mathbb{E}e^{-it\cdot X_{0}}\right|$
$=\left|\mathbb{E}\cos(t\cdot(X_{i}-X_{0}))+i\mathbb{E}\sin(t(X_{i}-X_{0}))\right.$
$\quad\left.-\left(\mathbb{E}\cos(t\cdot X_{i})+i\mathbb{E}\sin(t\cdot X_{i})\right)\left(\mathbb{E}\cos(t\cdot X_{0})-i\mathbb{E}\sin(t\cdot X_{0})\right)\right|$
Using the relations $\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y),$ $\cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y)$, $\sin(x)=-\sin(-x)$ and $\cos(x)=\cos(-x)$, the latter expression can be estimated by $I_{1}+I_{2}+I_{3}+I_{4}$, where
$I_{1}$ $:=|\mathrm{cov}(\cos(t\cdot X_{i}),\cos(t\cdot X_{0}))|,$
$I_{2}$ $:=|\mathrm{cov}(\sin(t\cdot X_{i}),\sin(t\cdot X_{0}))|,$
$I_{3}$ $:=|\mathrm{cov}(\sin(t\cdot X_{i}),\cos(t\cdot X_{0}))|,$
$I_{4}$ $:=|\mathrm{cov}(\cos(t\cdot X_{i}),\sin(t\cdot X_{0}))|.$
For $I_{2},I_{3},I_{4}$, we note that
$\max\{\mathbb{E}|\sin(t\cdot X_{i})|^{r},\mathbb{E}|\sin(t\cdot X_{0})|^{r}\}\leq|t|^{r}\mathbb{E}|X_{0}|^{r}<\infty.$
On the other hand, $\max\{\mathbb{E}|\cos(t\cdot X_{i})|,\mathbb{E}|\cos(t\cdot X_{0})|\}\leq 1$. Using the covariance inequality (2.7) for strong mixing random fields,
$\max\{I_{2},I_{3},I_{4}\}\leq 6|t|\|X_{0}\|_{r}\alpha^{1/s}(|i|).$
At last, having $\cos(x)=1-2\sin^{2}(x/2)$ and $\sin^{2}(x)\leq|\sin(x)|\leq|x|$, we arrive at
$I_{1}$ $=|\mathrm{cov}(\cos(t\cdot X_{i}),1-2\sin^{2}(t\cdot X_{0}/2))|$
$=2|\mathrm{cov}(\cos(t\cdot X_{i}),\sin^{2}(t\cdot X_{0}/2))|$
$\leq 6|t|\|X\|_{r}\alpha^{1/s}(|i|).$
For $|t|>1$, we simply use the covariance inequality (2.8) for bounded random variables $X=e^{it\cdot X_{i}},Y=e^{-it\cdot X_{0}}$ to write
$\left|\varphi_{X_{i}-X_{0}}(t)-|\varphi_{X_{0}}(t)|^{2}\right|=|\mathrm{cov}(Y,X)|\leq 4\alpha(|i|).$
The proof of Theorem 2 is analogical. In the proof of Lemma 1, we would use two times the covariance inequality (2.8) to arrive at the respective estimate
$\left|\varphi_{X_{i}-X_{0}}(t)-|\varphi_{X_{0}}(t)|^{2}\right|$ $\leq c\min\{\alpha(|i|)|t|\|X_{0}\|_{\infty},\alpha(|i|)\}\quad\text{for all }t\in\mathbb{R}^{d}.$
Proof.
(of Corrolary 2) Assume first that $p_{0}$ is centered. It is a known fact (see Section 2.1.1 of [6]) that if $\mathrm{cov}(p_{0},p_{i})=O(|i|^{-\gamma})$ for some $\gamma>d$ and if the spectral density $f(\lambda)=\sum_{j\in\mathbb{Z}^{d}}\mathbb{E}p_{0}\cdot p_{j}e^{ij\cdot\lambda}$ is bounded from below by a positive constant on $[0,2\pi]^{d}$, then
$\alpha(n)=O(n^{d-\gamma}).$
Keeping $\gamma>2d$ allows one to apply Theorem 1 with any $1<s<\frac{\gamma-d}{d}$. Finally, the statement of the corollary remains true for non-centered Gaussian perturbations, since the transformation $\Xi\mapsto\Xi+\mu=\{i+p_{i}+\mu+U\}$ has no impact on the number variance.
3 Gaussian perturbations
If we further assume Gaussianity of the perturbation, it is possible to explicitly determine the K-function, and hence also the Besag L-function.
To describe the joint distribution of $(p_{i},p_{j})$ for two perturbations $p_{i}$ and $p_{j}$, we need to introduce some matrix formalism. First, we will use the notation $\mathbf{I}_{n}$ for the $n\times n$ unit matrix. Moreover, $\mathbf{0}_{n\times p}$ denotes the $n\times p$ matrix of zeros. For an $n\times n$ matrix $A$, $\mathrm{tr}(A)$ is the trace of $A$ and for a general $n\times p$ matrix $B$, we let $\mathrm{vec}(B)$ be the vectorization of $B$. Finally, $A\otimes B$ denotes the tensor product of two matrices $A$ and $B$.
To define stationary Gaussian random perturbation field on $\mathbb{R}^{d}$, we assume that there is a positive definite matrix $\mathbf{V}\in\mathbb{R}^{d\times d}$ and for any $n\in\mathbb{N}$ and $(i_{i},\ldots i_{n})^{T}\in\mathbb{Z}^{n\times d}$, there is a matrix $\mathbf{M}$ with common rows equal to some $\mu\in\mathbb{R}^{d}$ and a positive definite matrix $\mathbf{U}\in\mathbb{R}^{n\times n}$ invariant w.r.t $\mathbb{Z}^{d}$-shifts of $(i_{1},\ldots,i_{n})$ such that the vector $(p_{i_{1}},\ldots,p_{i_{n}})$ has a $n\times d$ dimensional matrix Gaussian distribution ${\mathcal{MN}}_{n,d}(\mathbf{M},\mathbf{U},\mathbf{V})$ given by the probability density function
$p(\mathbf{X})=\frac{\exp\left(-\frac{1}{2}\mathrm{tr}\left[\mathbf{V}^{-1}(\mathbf{X}-\mathbf{M})^{T}\mathbf{U}^{-1}(\mathbf{X}-\mathbf{M})\right]\right)}{(2\pi)^{nd/2}|\mathbf{V}|^{n/2}|\mathbf{U}|^{d/2}},\quad\mathbf{X}\in\mathbb{R}^{n\times d}.$
Here, the matrix $\mathbf{U}$ determines the covariance structure among the perturbations while $\mathbf{V}$ respects the covariance structure within an individual perturbation. Specifically, $\mathbf{p}$ consists of independent Gaussian perturbations if $\mathbf{U}$ is a diagonal matrix.
Moreover, we call the Gaussian perturbation field regular if for any $n\in\mathbb{N}$ and $(i_{i},\ldots i_{n})^{T}\in\mathbb{Z}^{n\times d}$, we have $\mathbf{V}=\mathbf{I}_{d}$ and $\mathbf{U}$ is symmetric with the diagonal points equal to $\sigma^{2}>0$. The corresponding perturbed lattice is called regular Gaussian perturbed lattice.
Example 1.
Let $\mathbf{p}$ be a regular Gaussian perturbation field. Specifically, we take $n=2$ and $i,j\in\mathbb{Z}^{d}$. Then
$\mathbf{U}=\begin{pmatrix}\sigma^{2}&\sigma_{ij}\\ \sigma_{ij}&\sigma^{2}\end{pmatrix}$
for some $\sigma_{i,i}\in\mathbb{R}$ and $\sigma^{2}>0$. Using the vectorization notation, we have that
$\mathrm{vec}(p_{i},p_{j})\sim\mathcal{N}_{2d}(\mathrm{vec}(\mathbf{M}),\mathbf{V}\otimes\mathbf{U})$
with
$\mathbf{V}\otimes\mathbf{U}=\begin{pmatrix}\mathbf{U}&\mathbf{0}_{2\times 2}&\cdots&\mathbf{0}_{2\times 2}\\ \mathbf{0}_{2\times 2}&\mathbf{U}&\cdots&\mathbf{0}_{2\times 2}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}&\ldots&\mathbf{U}\end{pmatrix}.$
This means that
$p_{i}\sim\mathcal{N}_{d}(\mu,\sigma^{2}\mathbf{I}_{d})$
and
$\mathrm{cov}(p_{i},p_{j})=\mathbb{E}p_{i}\cdot p_{j}=d\sigma_{ij}.$ (3.1)
Proposition 1.
Assume that $\Xi$ is a regular perturbed Gaussian lattice with $\sigma^{2}>0$ and $\sigma_{i}:=d^{-1}\mathrm{cov}(p_{i},p_{0})<\sigma^{2}$ for all $i\in\mathbb{Z}^{d}\setminus\{0\}.$ The $K$-function of $\Xi$ is
$K(r)=\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}P_{d}(r^{2}/\kappa_{i},|i|^{2}/\kappa_{i}),\quad r\geq 0,$ (3.2)
where $\kappa_{i}:=2\sigma^{2}-2\sigma_{i}$ and $P_{d}(\cdot,\lambda)$ denotes the cumulative distribution function of the non-central chi-squared distribution with $d$ degrees of freedom and non-centrality parameter $\lambda$.
Proof.
Since the intensity of the perturbed lattice is $1$, we may write
$K(r)$ $=\alpha^{(2)}_{red}(B(0,r))$
$=\mathbb{E}\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbf{1}{\{i+p_{i}-p_{0}\in B(0,r)\}}$
$=\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbb{P}(i+p_{i}-p_{0}\in B(0,r))$
$=\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbb{P}(|i+p_{i}-p_{0}|\leq r)$
where the third equality is justified by the non-negativity of the summands. Let us now determine the distribution of $p_{i}-p_{0}$. By assumptions, the matrix $\mathbf{X}=(p_{0},p_{i})^{T}$ has matrix normal distribution $\mathcal{MN}_{2,3}((\mu,\mu)^{T},\mathbf{U},\mathbf{V})$ with
$\mathbf{U}=\begin{pmatrix}\sigma^{2}&\sigma_{i}\\ \sigma_{i}&\sigma^{2}\end{pmatrix},\quad\mathbf{V}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
We apply a linear transformation $Y_{i}:=\mathbf{D}\mathbf{X}\mathbf{C}$ with $\mathbf{D}=\begin{pmatrix}-1&1\end{pmatrix}$ and $\mathbf{C}=\mathbf{I}_{d}$. Then $Y_{i}=p_{i}-p_{0}$ has matrix normal distribution $\mathcal{MN}_{1,d}(\mathbf{D}\mathbf{M}\mathbf{C},\mathbf{D}\mathbf{U}\mathbf{D}^{T},\mathbf{C}^{T}\mathbf{V},\mathbf{C})$ $=\mathcal{MN}_{1,d}(\mathbf{0}_{1\times d},2\sigma^{2}-2\sigma_{i},\mathbf{I}_{d})=\mathcal{N}_{d}(\mathbf{0}_{1\times d},\kappa_{i}\mathbf{I}_{d})$. At last, the random vector $Z_{i}:=\kappa_{i}^{-1/2}(i+Y_{i})$ has a distribution $\mathcal{N}_{d}(\kappa_{i}^{-1/2}i,\mathbf{I}_{d})$ and we may write
$K(r)$ $=\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbb{P}(|i+Y_{i}|\leq r)$
$=\sum_{i\in\mathbb{Z}^{d}\setminus\{0\}}\mathbb{P}(|Z_{i}|^{2}\leq\kappa_{i}^{-1}r^{2}).$
Clearly $|Z_{i}|^{2}$ has a non-central chi-squared distribution with $d$ degrees of freedom and non-centrality parameter $\lambda=\kappa_{i}^{-1}|i|^{2}.$
Remark.
In case $\sigma^{2}>0$, $\kappa_{i}=0$ corresponds to the perfect correlation between the entries $p^{k}_{0}$ and $p^{k}_{i}$, $k=1,\ldots,d$ and hence between $p_{0}$ and $p_{i}$. In such a case, $p_{i}-p_{0}$ almost surely equals $\mathbf{0}_{1\times d}$. Let $\mathcal{A}$ be the subset of $\mathbb{Z}^{d}$ of all indexes $i$ such that $\kappa_{i}=0$. Then
$K(r)=|\mathcal{A}\cap B(0,r)|+\sum_{i\in\mathbb{Z}^{d}\setminus\{\mathcal{A},0\}}P_{d}(r^{2}/\kappa_{i},|i|^{2}/\kappa_{i}),\quad r\geq 0$
However, due to the assumption strict stationarity, we also have that $p_{2i}-p_{i}=\mathbf{0}_{1\times d}$ a.s. From this, if $\sigma_{i}=\sigma^{2}$ for some $i\in\mathbb{Z}^{d}$, then also $\sigma_{ki}=\sigma^{2}$ for any $k\in\mathbb{N}$ which clearly contradicts the assumptions of Corollary 2. To see that, we show that $p_{ki}-p_{0}=\mathbf{0}_{1\times d}$ a.s. for any $k\in\mathbb{N}$. Without loss of generality, we take $k=2$. Then
$\mathrm{cov}(p_{2i},p_{0})=\mathrm{cov}(p_{2i}-p_{i}+p_{i},p_{0})=\mathrm{cov}(p_{i},p_{0})=d\sigma^{2}.$
On the other hand, $\sigma^{2}=0$ produces simply a so called stationarized lattice $\mathbb{Z}^{d}+U$ for which the $K$-function has a trivial expression
$K(r)=|\mathbb{Z}^{d}\setminus\{0\}\cap B(0,r)|,\quad r\geq 0.$
Remark.
The terms in the sum (3.2) decay exponentially fast if $|i|-r$ increases with $i$. For practical simulation reasons, we are interested only in $r\in[0,b]$ for some fixed (usually not so big) $b>0$. It is therefore possible to approximate the $K$-function by numerically stable estimate
$\hat{K}(r)=\sum_{i\in\mathbb{Z}^{d};|i|\in[r-q,r+q]}P_{d}(r^{2}/\kappa_{i},|i|^{2}/\kappa_{i}),\quad r\in[0,b]$
for some suitable $q>0$ which depends on the dimension.
(a) theoretical $K$-function
(b) theoretical centered $L$-function
Figure 1: Theoretical values of the $K$-function and the centered $L$-function of the lattice perturbed by independent Gaussian random variables with standard deviation $0.25$.
4 Real data analysis
4.1 Description of the data
Nickel titanium, also known as nitinol (or NiTi), is a metal alloy of nickel and titanium. Nitinol alloys are widely used in medicine, battery development, space engineering, or other fields that benefit from their recoverable strain, which can be repeatedly induced by thermal and/or mechanical actuation.
In this paper, the grain microstructure in a polycrystalline NiTi wire was reconstructed from an experimental 3D synchrotron x-ray diffraction (3D-XRD) data set. The 3D-XRD technology provides information on microstructure geometry, especially the grain center position, grain volumes, and grain orientations. However, it does not provide the full geometry of the grains, such as its exact boundaries. That is a subject of further numerical studies based on a well-fitted mathematical models. For our study, we have a specimen at our disposal consisting of positions of 8063 grain centers which are contained in a 3D cylindric observation window of height 80 micrometers and radius $40$ micrometers. This data set was collected by the authors in [20] using the so-called cross-entropy method applied to 3D X-ray diffraction measurements. To eliminate border effects in our study, we extract a 3D cubic segment of size 35x35x70 micrometers that includes 4807 grain centers.
Figure 2: Positions of grain centers in NiTi alloy extracted using 3D-XRD.
From the perspective of model fitting, this data set was extensively studied in [22, 20]. We stick with the premise of homogenity of the data that was tested in the aforementioned papers. The plots of the estimated characteristic functions (see Figure 3 below) indicate that there is rather repulsive behavior between the points in the data set.
(a) $\hat{L}(r)-r$
(b) $\hat{g}$
(c) $\hat{G}$
Figure 3: Estimated (a) centered L-function, (b) pair-correlation function and (c) G-function. The estimated values are plotted in solid lines, resp. red dashed line, while being compared to a theoretical values of homogeneous Poisson point process (green dashed line).
4.2 Relation to the existing literature
Primarily, we compare our approach with the methodology proposed in [22]. To model the positions of the grain centers, the authors construct so called repulsive energy model, which follows the standard Gibbs formalism having finite volume density with respect to unit rate Poisson point process
$p_{k}(\mathbf{y})=\frac{1}{Z}\beta^{\#(\mathbf{y})}\prod_{i=1}^{k-1}\gamma_{i}^{\sum_{y_{1},y_{2}\in\mathbf{y}}^{\neq}\mathbf{1}\{\|y_{1}-y_{2}\|\in(\delta_{i-1},\delta_{i}]\}},\quad\mathbf{y}\in\mathbf{N}_{f},$ (4.1)
where $k\in\mathbb{N}$ is an apriori given parameter, $Z$ is the normalizing constant and $\beta>0,0\leq\gamma_{1},\ldots,\gamma_{k-1}\leq 1$ and $0=\delta_{0}<\delta_{1}<\ldots<\delta_{k-1}$ are unknown parameters of the model.
Lemma 2.
A point process given by finite dimensional densities (4.1) w.r.t the unit rate Poisson process is not hyperuniform provided that $\delta_{k-1}<\infty$.
Proof.
In the setting of Section 4.1 in [5], the density (4.1) corresponds to the pairwise energy function
$H(\mathbf{y})=-\sum_{y_{1},y_{2}\in\mathbf{x}}^{\neq}\sum_{i=1}^{k-1}{\mathbf{1}\{\|y_{1}-y_{2}\|\in(\delta_{i-1},\delta_{i}]\}}\log\gamma_{i}.$
Due to the condition posed on the parameters $\gamma_{1},\ldots,\gamma_{k-1}$, the energy $H$ is non-negative regardless the parameter $k$ and therefore, satisfies both superstability and lower-regularity assumptions. Moreover, $\mathbf{y}\in\mathbf{N}_{f}$ and $x\in\mathbb{R}^{d}$, the Papangelou conditional intensity is given by
$\lambda_{k}^{*}(x,\mathbf{y})=\frac{p_{k}(\mathbf{y}\cup\{x\})}{p_{k}(\mathbf{y})}=\beta\prod_{i=1}^{k-1}\gamma_{i}^{\sum_{y\in\mathbf{y}}\mathbf{1}\{\|y-x\|\in(\delta_{i-1},\delta_{i}]\}}.$
Define a function $g_{k}:\mathbb{R}^{d}\to\mathbb{R}$ by
$g_{k}(x)=\frac{\lambda_{k}^{*}(0,\mathbf{y}\cup\{x\})}{\lambda_{k}^{*}(0,\mathbf{y})}=\prod_{i=1}^{k-1}\gamma_{i}^{\mathbf{1}\{\|x\|\in(\delta_{i-1},\delta_{i}]\}}.$
The function $g_{k}$ does not depend on the choice of $\mathbf{y}\in\mathbf{N}_{f}$. With the convention $10^{0}=$ 1,
$\int_{\mathbb{R}^{d}}\left|1-g_{k}(x)\right|\mathrm{d}\,x=\int_{\cup_{i=1}^{k-1}(\delta_{i-1},\delta_{i}]}\left|1-g_{k}(x)\right|\mathrm{d}\,x\leq\sum_{i=1}^{k-1}\delta_{i}-\delta_{i-1}=\delta_{k-1}<\infty.$
The resulting process is therefore not hyperuniform by Corollary 3 of [5].
Note that this statement is in accordance with the fact that the choice $k=1$ corresponds directly to the Poisson point process with intensity $\beta$ and $k=2$ defines the Strauss process. The authors suggest modeling the data with $k=3$. That requires estimating $5$ unknown parameters of the model: $\beta,\delta_{1},\delta_{2},\gamma_{1},\gamma_{2}$. In the presented methodology, we restrict ourselves to a maximum of $3$ unknown parameters. This allows us to perform faster computations compared to the latter approach.
4.3 Hyperuniformity in the NiTi data
Normalizing the intensity. The data set was first rescaled to unit intensity. The total number of points 4807 distributed in a box of volume $70^{3}\mu m^{3}$ were resized into a box $W:=[-8.4384,8.4384]^{3}$ to gain unit intensity and the center of the box being at the origin. That is in order to get us in the context of the paper [16] that will be used to estimate the hyperuniformity exponent and also to have the same intensity as our proposed model, a perturbed lattice. However, the latter is not necessary because the perturbed lattices can easily be defined to have any desired intensity.
Hyperuniformity. It is far from being true that repulsive point patterns need to yield hyperuniformity, note for instance the Matérn III hard-core process (also known under RSA model) or short-range hard-core Gibbs models (see [5]). A very brief, yet only empirical approach to asses hyperuniformity is visualizing the tails of the distribution of the number of points using a simple histogram. In Figure 5 below, we extracted 64 disjoint boxes such that their distance is $1.5$. This is the distance at which the joint correlation between two points becomes negligible, as seen in the estimation of the pair correlation function of the rescaled data set (Figure 4). We may observe that the realization of the Poisson point process has slightly heavier tails, which could be understood as a preliminary sign of the potential presence of hyperuniformity.
Figure 4: Estimated pair correlation function of the rescalled data set. After $r=1.5$, it becomes negligible.
(a) NiTi data
(b) Poisson point process
Figure 5: Histograms presenting the number of points appearing in 64 distant boxes of the size of $2.7^{3}$.
Statistical inference for hyperuniformity has gained attention only recently, in particular in [8, 12] and [16]. The first one assumes having more realizations of the observed data, which is not the case in our paper. The latter two tests work efficiently with only one sample. However, the sensibility goes firmly in hand with the size of the data set. In dimension $d=2$, the recommendations vary around at least 5000 points. It is not specified for higher dimensions, but the necessity of a large sample grows proportionally with dimension.
Following [16], we performed a test for the hyperuniformity exponent (see Eq. (2.5)) for our data, resulting in $\hat{\alpha}=0,81$ speaking in favor of hyperuniformity. The test consists of carefully choosing suitable parameters and test Schwartz functions (tapers) and we provide detailed analysis in the R script at https://github.com/DanielaFlimmel/Fitting_PL. In conclusion, it is reasonable to assume that the data come from a hyperuniform model.
4.4 Models description
We assume three different models, among which two Gaussian perturbed lattices, for which we may use the minimum contrast method together with Proposition 1 for estimating the respective parameters. Throughout this section, $U$ always stands for the uniform random variable in $[0,1]^{d}$ that is assumed to be independent of any other random variable mentioned in the following.
Independent Gaussian perturbations. In physical papers, this model is also called an Einstein pattern. Assume a one-parametric model of the Gaussian perturbed lattice
$\Xi_{iid}^{\sigma}=\{i+U+p^{\sigma}_{i},i\in\mathbb{Z}^{3}\},$
where $\mathbf{p}=\{p^{\sigma}_{i},i\in\mathbb{Z}^{d}\}$ are independent, identically distributed centered Gaussian random variables in $\mathbb{R}^{3}$ with variance matrix $\mathbf{V}=\sigma^{2}\mathbf{I}_{3}$. Clearly, $\Xi_{iid}$ forms a hyperuniform point process.
Dependent Gaussian perturbations. We assume a three-parametric model of dependently Gaussian perturbed lattice
$\Xi_{exp}^{\sigma,r,\gamma}=\{i+U+p_{i}^{\sigma,r,\gamma};i\in\mathbb{Z}^{3}\}.$
where $\mathbf{p}=\{p_{i}^{\sigma,r,\gamma};i\in\mathbb{Z}^{3}\}$ is Gaussian random field with exponentially decaying covariances defined as follows: As in Example 1, we set the variance matrix $\mathbf{V}=I_{3}$ and for $n=2$, we prescribe the covariance matrix between $p_{i_{1}},p_{i_{2}}$ by
$\mathbf{U}=\begin{pmatrix}\sigma^{2}&\sigma^{2}\exp\{-|i_{1}-i_{2}|^{\gamma}/r\}\\ \sigma^{2}\exp\{-|i_{1}-i_{2}|^{\gamma}/r\}&\sigma^{2}\end{pmatrix},$
where $\sigma>0,r>0$ and $\alpha\in[0,2]$ are subjects to model fitting. The restriction on $\alpha$ comes from the necessity of positive definiteness of the matrix $\mathbf{V}\otimes\mathbf{U}$. Since
$\sum_{j\in\mathbb{Z}^{3}}|\mathrm{cov}(p_{0}^{\sigma,r,\gamma},p_{i}^{\sigma,r,\gamma})|=3\sigma^{2}\sum_{j\in\mathbb{Z}^{d}}\exp\{-|j|^{\gamma}/r\}<\infty,$
the spectral density exists and is positive by Theorem 1.6. of [1]. The assumptions of 2 are satisfied and therefore, the point process $\Xi_{exp}^{\sigma,r,\alpha}$ is hyperuniform.
Perturbation of cloaked lattices. The concept of cloaked and perturbed lattices was introduced in [11]. As suggested in [16] in Section 4.2.1, we start with the cloaked lattice
$\Xi_{c}:=\{i+U_{i}+U,i\in\mathbb{Z}^{d}\},$
where $U$ and $\{U_{i},i\in\mathbb{Z}^{3}\}$ are i.i.d uniform random variables on $[0,1]^{3}$. Next, we perturb the cloaked lattice by a perturbation field $\{p_{i},i\in\mathbb{Z}^{3}\}$, where $p_{i}$ are i.i.d random variables with characteristic function $\phi$ satisfying
$1-|\phi(k)|^{2}\sim c|k|^{{\alpha}}\quad\text{as }|k|\to 0$ (4.2)
for some $c>0$. In practice, it is recommended to use $\alpha=\hat{\alpha}$, the estimated hyperuniformity exponent from the data. This model forms a hyperuniform point process with the target exponent $\alpha$ of the structure factor in (2.5). To obtain the order (4.2), we choose
$p_{0}\stackrel{{\scriptstyle\mathcal{D}}}{{=}}\sqrt{Y(\alpha/2)}N,$ (4.3)
where $Y(\alpha/2)$ is the one-sided $\alpha/2$-stable law and $N\sim N_{3}(0,\sigma^{2})$. To simulate from the distribution of $Y$, we use Corollary 4.1 of [10] Eventually, fitting this model is a subject of estimating the single parameter $\sigma>0$.
4.5 Model fitting
The implementation of the models together with simulation procedures is provided in the R code which is available at https://github.com/DanielaFlimmel/Fitting_PL. To eliminate the border effects of the simulations, we simulated the perturbed lattices in extended windows and then cropped it to match the size of the data.
Minimal contrast method. In the case of Gaussian perturbations, we have Lemma 1 in hand, and we can use the minimal contrast method directly for estimating the parameters of the models. This means to minimize
$D(\theta)=\int_{r_{1}}^{r_{2}}\left|\left(\sum_{i\in\mathbb{Z}^{3}\setminus\{0\}}P_{3}(r^{2}/\kappa^{\theta}_{i},|i|^{2}/\kappa^{\theta}_{i})\right)^{1/4}-\hat{K}_{NiTi}(r)^{1/4}\right|^{2}\mathrm{d}\,r,$ (4.4)
where $\theta$ belongs to some parametric space given by the model and $\kappa_{i}^{\theta}$ depends on $\theta$. The estimated K-function of the NiTi data is $\hat{K}_{NiTi}$. For computational purposes, we used $\hat{K}(r)$ in (4.4) instead of the theoretical $K$-function with $q=15$ (see the remark after Proposition 1).
Independent Gaussian perturbations. For independent Gaussian perturbations, we have $\kappa_{i}^{\sigma}=2\sigma^{2}$. We choose $r_{1}=0,r_{2}=3$ and the final minimizer is $\hat{\sigma}=0.18$. Below is the comparison of the centered Besag’s L-function of the fitted model together with the estimated centered L-function from the data.
Figure 6: (a) Black lines corresponds to the estimated centered L-function, green dashed line is the theoretical value of Poisson point process while red line represents the theoretical value of the fitted independently perturbed lattice with Gaussian random variables with standard deviation $\sigma=0.18$. (b) Global envelope test for the fitted i.i.d perturbed Gaussian lattice.
The periodic peaks that occur at the tail of the theoretical centered $L$-function in Figure 6 are typical for independently perturbed lattices. Note that this phenomenon is less visible if we increase the variance in the model. However, the estimated standard deviation is low enough to preserve this property. To verify the validity of the model, we performed a global envelope test based on [18] (see Figure 6, (b)). Due to these tail peaks and also suppression of the variance caused by hyperuniformity of the model, the independently perturbed lattice by Gaussian random variables is not a good approximation of the data. We need a model where the peaks are less visible (cloaked). This is solved either if we assume dependencies among the perturbations or if we use a distribution for the perturbations with heavier tails.
Dependent Gaussian perturbations. Similarly as with the independent Gaussian perturbation, we used the minimum contrast method based on the $K$-function of Lemma 1 and (4.4) to estimate the parameters in the dependent Gaussian perturbation model resulting in $\hat{\sigma}=0.3,\hat{r}=2.5$ and $\hat{\gamma}=2$. See below in Figure 7 the comparison of the theoretical centered $L$-function of the fitted model and the estimated function from the data. The tail peaks peaks are less visible then in the i.i.d case. However, the global envelope test rejects this model. The reason is the low variance of the $L$-function in the region $r\in(0.5,1.5)$.
Figure 7: (a) Black lines correspond to the estimated centered L-function, green dashed line is the theoretical value of Poisson point process while red line represents the theoretical value of the fitted dependently perturbed lattice with parameters $\hat{\sigma}=0.3,\hat{r}=2.5$ and $\hat{\gamma}=2$ (b) Global envelope test for the fitted dependent perturbed Gaussian lattice.
For this reason, we performed the second step of the optimization procedure. Based on previous simulations, we compute the empirical variance of the centered $L$-function for each $r\in[0,3]$. Let us denote this function by $\hat{l}(r)$. The second optimization procedure is then based on a weighted function
$D(\theta)=\int_{r_{1}}^{r_{2}}\frac{1}{\sqrt{\hat{l}(r)}}\left|\left(\sum_{i\in\mathbb{Z}^{3}\setminus\{0\}}P_{3}(r^{2}/\kappa^{\theta}_{i},|i|^{2}/\kappa^{\theta}_{i})\right)^{1/4}-\hat{K}_{NiTi}(r)^{1/4}\right|^{2}\mathrm{d}\,r,$
With $r_{1}=0.2$ and $r_{2}=2$, we arrive at the estimates $\hat{\sigma}=0.32,\hat{r}=2.5$ and $\hat{\gamma}=2$. Indeed, this model performed better compared to the first attempt.
Figure 8: (a) Black lines corresponds to the estimated centered L-function, green dashed line is the theoretical value of Poisson point process while red line represents the theoretical value of the fitted dependently perturbed lattice with parameters $\hat{\sigma}=0.32,\hat{r}=2.5$ and $\hat{\gamma}=2$ (b) Global envelope test for the fitted dependent perturbed Gaussian lattice. The corresponding p-value obtained by the global area rank envelope test is $24,4\%$
Perturbation of cloaked lattices. The one-parametric model with independently perturbed cloaked lattice by random variables with distribution given by (4.3) was implemented. However, a preliminary analysis of the realizations for different choices of the parameter $\sigma$ of the model showed that this is not a suitable option for our data. First, for the standard deviation being in the region $\sigma\in(0,0.25)$, the peaks inherited from the lattice are too visible in the centered $L$-function. On the other hand, for $\sigma>0.25$, the centered $L$-function starts to look Poisson-like. This is demonstrated in Figure 9 where we simulated 50 realizations for different choices of $\sigma$.
(a) $\sigma=0.1$
(b) $\sigma=0.25$
(c) $\sigma=0.5$
Figure 9: The centered $L$-function estimated from the data (black) compared with 50 realizations of the perturbed cloaked lattice (grey) with (a) $\sigma=0.1$ (b) $\sigma=0.25$ and (c) $\sigma=0.5$.
5 Conclusion
When analyzing a data set that exhibits some degree of regularity, it is important to examine how quickly the number variance grows in comparison to a Poisson point process. If there is a statistical or physical justification for hyperuniformity, a hyperuniform point process may serve as a suitable model for the data. In particular, perturbed lattices offer a computationally efficient and easy-to-implement option. Specifically, independently perturbed lattices share these advantages, although they provide only a rough approximation due to the visible Bragg peaks. For a more precise representation of the relationships among points, we recommend incorporating dependencies among the perturbations. Especially, dependent Gaussian perturbations have proven effective, as their $K$-function is explicitly known.
Using perturbed lattices offers a numerical advantage over the traditional Gibbsian approach. However, Gibbs point processes should still be considered for regular data. The energy functions in Gibbs models provide a deeper understanding and greater control over the relationship among the points, making them still a slightly more flexible modeling choice than perturbed lattices.
6 Future extension
Central limit theorems for cell characteristics of a tessellation based on the NiTi data. First, we should find a suitable model for the distribution of the radii $r$ (e.g. by following the hierarchical approach by [22]) to construct a marked perturbed lattice $\Xi_{m}=\{(x,m),x\in\Xi,r\geq 0\}=\{(x_{i},r_{i})_{i\in\mathbb{Z}^{3}}\}$, where the cell generated by a marked point $(x,r)$ is
$C^{\rho}((x,r),\Xi_{m})=\bigcap_{i\in\mathbb{Z}^{3}}\{y\in\mathbb{R}^{3}:\rho(y,(x,r))\leq\rho(y,(x_{i},r_{i}))\}.$
The common choices of the function $\rho$ are listed below
  • Voronoi tessellation: $\rho(y,(x,r))=\|y-x\|$
  • Laguerre tessellation: $\rho(y,(x,r))=\|y-x\|^{2}-r^{2}$
  • Johnson-Mehl tessellation: $\rho(y,(x,r))=\|y-x\|-r$
If $K_{\mathbf{0}}^{\rho}(\Xi_{m})$ is the cell generated by the typical point of $\Xi_{m}$ (in the Palm sense), we are interested in the estimation of $\mathbb{E}\xi(K_{\mathbf{0}}^{\rho})$ where $\xi$ is some function on closed sets (diameter, volume, etc.)
$H_{n}=\sum_{i\in\mathbb{Z}^{3}\cap W_{n}}\xi(C((x_{i},r_{i}),\Xi_{m}))\cdot\text{weight}_{i},$
where the weights are chosen reasonably so that the estimator $H_{n}$ is (asymptotically) unbiased and consistent. The stabilization methods and mixing properties of the perturbation field could then be used to show that
$\frac{H_{n}-\mathbb{E}H_{n}}{\sqrt{\sigma_{n}}}\xrightarrow[n\to\infty]{\mathcal{D}}N(0,\sigma^{2}_{\xi}),$
where $\sigma_{n}=\textbf{var}|\Xi\cap W_{n}|$. This asymptotic study should be carried out in a separate paper.
Acknowledgement
This work was supported by the Czech Science Foundation, project no. 22-15763S.
I am thankful to Gabriel Mastrilli for his help in implementing the test for the hyperuniformity exponent in dimension $d=3$ and also Jiří Dvořák for his helpful recommendations.
Special thanks go to my husband and my parents in law who supported my research by helping me with the maternal duties.